# Aligning Text to Image in Diffusion Models is Easier Than You Think

### v3

@@ Abstract
While recent advancements in generative modeling have significantly improved text-image alignment, some residual misalignment between text and image representations still remains. Although many approaches have attempted to address this issue by fine-tuning models using various reward models, etc., we revisit the challenge from the perspective of representation alignment—an approach that has gained popularity with the success of REPresentation Alignment (REPA) [36]. We first argue that conventional text-to-image (T2I) diffusion models, typically trained on paired image and text data (i.e., positive pairs) by minimizing score matching or flow matching losses, is suboptimal from the standpoint of representation alignment. Instead, a better alignment can be achieved through contrastive learning that leverages both positive and negative pairs. To achieve this efficiently even with pretrained models, we introduce a lightweight contrastive fine-tuning strategy called SoftREPA that uses soft text tokens. This approach improves alignment with minimal computational overhead by adding fewer than 1M trainable parameters to the pretrained model. Our theoretical analysis demonstrates that our method explicitly increases the mutual information between text and image representations, leading to enhanced semantic consistency. Experimental results across text-to-image generation and text-guided image editing tasks validate the effectiveness of our approach in improving the semantic consistency of T2I generative models. Project Page: https://softrepa.github.io/
Figure 1: Representative results for image generation and image editing. SoftREPA provides much improved text to image alignment by introducing a negligible size of learnable soft tokens.
@@ 1 Introduction
Achieving effective alignment between modalities is essential in multimodal generative modeling. Specifically, latent diffusion models [26] enable various conditioning mechanisms, such as text, during the generation process for the semantic alignment between the two modalities.
For UNet-based denoisers [26, 22, 24], image representations are updated to align with fixed text representations encoded by pre-trained CLIP models through cross-attention layers. In contrast, transformer-based denoisers [23, 7] jointly update both text and image representations by concatenating them and processing them through self-attention layers. However, there is still room for further improvement, and to mitigate the remaining misalignment between text and image representations, a more effective representation learning approach is necessary.
Recent advancements in representation alignment for Diffusion Transformers (DiT) [23] have significantly enhanced their ability to learn semantically meaningful internal representations [36, 9]. Notably, REPA [36] demonstrated that aligning the internal representations of DiT with an external pre-trained visual encoder during training significantly improves both discriminative and generative performance.
Figure 2: Network architecture and algorithmic concept of SoftREPA. (a) Learnable soft tokens are prepended to the text features across the upper layers, with each layer’s previous soft tokens being replaced by the soft tokens of the current layer. (b) The soft tokens are optimized to contrastively match the score with positively conditioned predicted noise while repelling the score from negatively conditioned predicted noise. This process implicitly sharpens the joint probability distribution of images and text by reducing the log probability of negatively paired conditions.
To enhance the textual semantic alignment of DiT’s internal representations, we leverage these ideas to further enhance text-image alignment in text-to-image (T2I) generative models. Specifically, we are interested in the contrastive learning framework, a widely used strategy for multimodal alignment in most vision-language representation learning approaches such as CLIP [25], ALBEF [15], BLIP [16], ALIGN [10].
@@
One of the most important contributions of this work is therefore the introduction of soft text tokens which are trained through contrastive text-and-image training. This allows the model to dynamically adapt its text representations, improving alignment with generated images without requiring full model finetuning. This simple yet effective approach significantly enhances text-to-image alignment in text-to-image generation and text-guided image editing tasks. Furthermore, the method is so flexible that it can be used with any pretrained T2I generative models. Theoretically, we show that this approach explicitly increases the mutual information between text and images, leading to better semantic consistency in multi-modal representations. We therefore refer to our method as SoftREPA, short for Soft REPresentation Alignment. Our contributions can be summarized as follows.
@@ 2 Preliminaries
Flow Models. Suppose that we have access to samples from target distribution $X_{0}\sim q$ and source distribution $X_{1}\sim p$. The goal of flow model is to generate $X_{0}$ starting from $X_{1}$. Specifically, we define a velocity field $v_{t}({\bm{x}})$ of a flow $\psi_{t}({\bm{x}}):[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ that satisfies $\psi_{t}(X_{0})=X_{t}$ and $\psi_{1}(X_{0})=X_{1}$. Here, the $\psi_{t}$ is uniquely characterized by a flow ODE:
@@
where $v_{t}({\bm{x}}_{t}|{\bm{x}}_{0})$ is a conditional velocity field that defines a conditional flow $\psi_{t}({\bm{x}}_{t}|{\bm{x}}_{0})$ satisfying $\psi_{t}({\bm{x}}_{1}|{\bm{x}}_{0})={\bm{x}}_{t}$.
In particular, the linear conditional flow defines the flow as ${\bm{x}}_{t}=\psi_{t}({\bm{x}}_{1}|{\bm{x}}_{0})=(1-t){\bm{x}}_{0}+t{\bm{x}}_{1}$. Then, we can compute the conditional velocity field $v_{t}({\bm{x}}_{t}|{\bm{x}}_{0})=\dot{\psi}_{t}(\psi_{t}^{-1}({\bm{x}}_{t}|{\bm{x}}_{0})|{\bm{x}}_{0})={\bm{x}}_{1}-{\bm{x}}_{0}$ [18]. Thus, the conditional flow matching loss is defined as
@@
$d{\bm{x}}_{t}$ $=\left(\mathbb{E}\left[\frac{{\bm{x}}_{t}-(1-t){\bm{x}}_{0}}{t}|{\bm{x}}_{t}\right]-\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]\right)dt$ (7)
$=\frac{{\bm{x}}_{t}-\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]}{t}dt$
which is the probability-flow ODE (PF-ODE) of DDIM [28, 11] when ${\bm{x}}_{1}\sim p:=\mathcal{N}(0,{\mathbf{I}}_{d})$. In other words, flow models and score-based models can be used interchangeably.
2.1 Contrastive Representation Learning
Contrastive learning aims to maximize the similarity between semantically related text-image pairs while pushing apart unrelated pairs in a shared representation space [5, 25]. Formally, given a batch of N image-text pairs ${(\textbf{I}_{i},\textbf{T}_{i})}^{N}_{i=1}$, contrastive learning optimizes a contrastive loss function, typically a variant of the InfoNCE loss:
$\mathcal{L}_{\text{CLIP}}=-\frac{1}{N}\sum^{N}_{i=1}\log\frac{\exp(\text{sim}(\textbf{I}_{i},\textbf{T}_{i})/\tau)}{\sum^{N}_{j=1}\exp(\text{sim}(\textbf{I}_{i},\textbf{T}_{j})/\tau)}$ (8)
where $\text{sim}(\textbf{I}_{i},\textbf{T}_{i})$ is a similarity metric such as cosine similarity or inner product between image and text embeddings, and $\tau$ denotes a temperature parameter that controls the distribution sharpness. This contrastive loss encourages image and text representations to form a joint multimodal representation space, where aligned pairs are close together, and unaligned pairs are separated. Several fundamental models have successfully applied contrastive learning for multi-modal representation learning [25, 10, 15, 16] and self-supervised learning [5].
@@ 3.1 Contrastive T2I Alignment Loss
Furthermore, in SoftREPA we define $l({\bm{x}}^{(i)},{\bm{y}}^{(j)})$ as the logit from the variation of the denoising score matching loss for the case of T2I diffusion models [31].
$l({\bm{x}},{\bm{y}})=e^{-\mathbb{E}_{t,{\bm{\epsilon}}}[\left\lVert{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t,{\bm{y}})-{\bm{\epsilon}}\right\rVert^{2}/\tau(t)]}$ (10)
@@
Here, $\tau(t)$ represents time scheduling parameter. One might consider using the denoising score matching loss, $-\mathbb{E}_{t,{\bm{\epsilon}}}[\left\lVert{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t,{\bm{y}})-{\bm{\epsilon}}\right\rVert^{2}]$, as a similarity measure instead of relying on its logit value. However, our experiments revealed that this formulation can introduce instability during training due to the unbounded nature of the similarity measure. To address this, SoftREPA introduce an exponential function to constrain the logit values, effectively stabilizing the training process.
Intuitively, as illustrated in Fig. 2(b), training a diffusion model with this loss function would lead to predicted noise with matching text description align with the true noise. Conversely, when conditioned on a mismatched text, the predicted noise would deviate further from the true noise. It implies that the conditional probability distribution of the trained model gets sharpened, increasing the distinction between different text conditions, which is expected to enhance image-text alignment in the generated images.
@@ 3.2 The Soft Tokens
To distill the contrastive score matching objective above, we introduce a learnable soft token ${\bm{s}}_{\psi}$, which is the only trainable parameter while pretrained model remains frozen. In particular, as illustrated in Fig. 2, we introduce a set of learnable soft tokens that vary between layers and are indexed by time $t$. The soft token at layer $k$ is defined using an embedding function with the number of soft tokens as $m$:
$\mathbf{s}^{(k,t)}_{\psi}=\text{Embedding($k,t$)}\in\mathbb{R}^{m\times d}$ (12)
At the layer $k$, the text representation is updated by concatenating the time-indexed soft token with text hidden representation from the previous layer: $\hat{\mathbf{H}}^{(k-1,t)}_{\text{text}}=[\mathbf{s}^{(k,t)}_{\psi};\mathbf{H}^{(k-1,t)}_{\text{text}}]\in\mathbb{R}^{(m+n)\times d}$.
The soft token is prepended on text features across the layers, leading to a modified denoising function: $v_{\theta}({\bm{x}}_{t},t,{\bm{y}},{\bm{s}}_{\psi})$. By adopting the soft tokens, we can adjust the image and text representations to get better semantic alignment. Additionally, for computational efficiency, we approximate the expectation within $l({\bm{x}},{\bm{y}})$ using a single Monte Carlo sample, effectively replacing it as
$\tilde{l}({\bm{x}},{\bm{y}},{\bm{s}}_{\psi})=e^{-\|v_{\theta}({\bm{x}}_{t},t,{\bm{y}},{\bm{s}}_{\psi})-({\bm{\epsilon}}-{\bm{x}}_{0})\|^{2}/\tau(t)}$ (13)
@@
$\mathcal{L}_{\text{SoftREPA}}(\psi)=-\mathbb{E}_{\begin{subarray}{c}({\bm{x}},{\bm{y}})\sim p_{\text{data}}\\ t\sim U(0,1)\\ {\bm{\epsilon}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})\end{subarray}}\log\frac{\exp(\tilde{l}({\bm{x}},{\bm{y}},{\bm{s}}_{\psi}))}{\sum_{j}\exp(\tilde{l}({\bm{x}},{\bm{y}}^{(j)},{\bm{s}}_{\psi}))}$ (14)
Then the learnable token ${\bm{s}}_{\psi}$ is optimized as follows:
$\mathbf{s}^{*}_{\psi}=\arg\min_{\psi}\mathcal{L}_{\text{SoftREPA}}(\psi)$ (15)
During inference, the trained soft tokens are leveraged with fixed text representation from the text encoder across layers and timesteps. The detailed procedure during the image generation process based on MM-DiT architecture is described in Algorithm 1.
Algorithm 1 Image Generation with Soft Tokens in MM-DiT
1:Gaussian noise $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
2:Text $\mathbf{y}\sim p_{\text{data}}$
3:Soft token $\mathbf{s_{\psi}}\sim\text{Embedding}(k,t)$
4:Number of layers $N$, Threshold layer $L$
5:Time steps $\{t_{T},t_{T-1},...,t_{0}\}$
6:Initialize $\mathbf{H}_{\text{img}}^{(0,T)}\leftarrow\mathbf{z}$
7:Initialize $\mathbf{H}_{\text{text}}^{(0,T)}\leftarrow\text{TextEncoder}(\mathbf{y})$
8:$n=|\mathbf{H}_{\text{text}}^{(0,T)}|$
9:for $t$ in $\{t_{T},t_{T-1},...,t_{0}\}$ do
10:    for $l=1$ to $N$ do
11:         if $k\leq L$ then
12:             $\mathbf{s_{\psi}}^{(k,t)}\leftarrow\text{Embedding}(k,t)$
13:             $\hat{\mathbf{H}}_{\text{text}}^{(k-1,t)}\leftarrow[\mathbf{s_{\psi}}^{(k,t)};\mathbf{H}_{\text{text}}^{(k-1,t)}]$
14:         else
15:             $\hat{\mathbf{H}}_{\text{text}}^{(k-1,t)}\leftarrow\mathbf{H}_{\text{text}}^{(k-1,t)}$
16:         end if
17:         $\mathbf{H}_{\text{img}}^{(k,t)},\hat{\mathbf{H}}_{\text{text}}^{(k,t)}\leftarrow\text{Layer}_{l}(\mathbf{H}_{\text{img}}^{(k-1,t)},\hat{\mathbf{H}}_{\text{text}}^{(k-1,t)})$
18:         $\mathbf{H}_{\text{text}}^{(k,t)}\leftarrow\hat{\mathbf{H}}_{\text{text}}^{(k,t)}[-n:,:]$triangleright Drop soft tokens
19:    end for
20:end for
21:return $\mathbf{\hat{X}}=\text{Decoder}(\mathbf{H}_{\text{img}}^{(N,t_{0})})$
@@ 3.3 Relationship with Mutual Information
This formulation closely resembles a cross-entropy loss, where the conditional log-likelihood $\hat{l}({\bm{x}},{\bm{y}})$ can be interpreted as logits.
From the fact that mutual information is the expectation of the pointwise mutual information, we derive:
@@
where ${\bm{x}}^{(i)}$ represents the $i$-th sample in the data, $n$ denotes the number of samples in the data, and $D$ indicates the constant.
Notably, Eq. (20) is very similar in form to the contrastive objective in Eq. (9), where the logit is the negative of the diffusion loss. This implies that minimizing our contrastive learning objective is closely related to maximizing the mutual information between image-text pairs under the diffusion model.

### v4

@@ Abstract
While recent advancements in generative modeling have significantly improved text-image alignment, some residual misalignment between text and image representations still remains. Some approaches address this issue by fine-tuning models in terms of preference optimization, etc., which require tailored datasets. Orthogonal to these methods, we revisit the challenge from the perspective of representation alignment—an approach that has gained popularity with the success of REPresentation Alignment (REPA) [45]. We first argue that conventional text-to-image (T2I) diffusion models, typically trained on paired image and text data (i.e., positive pairs) by minimizing score matching or flow matching losses, is suboptimal from the standpoint of representation alignment. Instead, a better alignment can be achieved through contrastive learning that leverages existing dataset as both positive and negative pairs. To enable efficient alignment with pretrained models, we propose SoftREPA—a lightweight contrastive fine-tuning strategy that leverages soft text tokens for representation alignment. This approach improves alignment with minimal computational overhead by adding fewer than 1M trainable parameters to the pretrained model. Our theoretical analysis demonstrates that our method explicitly increases the mutual information between text and image representations, leading to enhanced semantic consistency. Experimental results across text-to-image generation and text-guided image editing tasks validate the effectiveness of our approach in improving the semantic consistency of T2I generative models. Project Page: https://softrepa.github.io/
Figure 1: Representative results for image generation and image editing. SoftREPA provides much improved text-to-image alignment by introducing a negligible size of learnable soft tokens.
@@ 1 Introduction
Achieving effective alignment between modalities is essential in multimodal generative modeling. Specifically, latent diffusion models [32] enable various conditioning mechanisms, such as text, during the generation process for the semantic alignment between the two modalities. For UNet-based denoisers [32, 28, 30], image representations are updated to align with fixed text representations encoded by pre-trained CLIP models through cross-attention layers. In contrast, transformer-based denoisers [29, 8] jointly update both text and image representations by concatenating them and processing them through self-attention layers. However, there is still room for further improvement, and to mitigate the remaining misalignment between text and image representations, a more effective representation learning approach is necessary.
Recent advancements in representation alignment for Diffusion Transformers (DiT) [29] have notably enhanced their ability to learn semantically meaningful internal representations [45, 11]. Notably, REPA [45] showed that aligning the internal representations of DiT with an external pre-trained visual encoder during training significantly improves both discriminative and generative performance. To further enhance text-image alignment in text-to-image (T2I) generative models, we leverage these ideas for aligning DiT’s internal representations.
@@
Specifically, we focus on the contrastive learning framework, a widely adopted strategy for multimodal alignment in vision-language representation learning, as seen in models such as CLIP [31], ALBEF [20], BLIP [21], and ALIGN [12]. One of the key contributions of this work is the introduction of soft text tokens, which are optimized via contrastive image-text training. These soft tokens allow the model to dynamically adapt its text representations, improving alignment with generated images without requiring full model fine-tuning. This simple yet effective approach significantly enhances text-to-image alignment in both text-to-image generation and text-guided image editing tasks. Notably, the method is highly flexible and can be seamlessly integrated with any pretrained text-to-image (T2I) generative model. From a theoretical standpoint, we show that this approach explicitly increases the mutual information between text and image representations, resulting in improved semantic consistency across modalities. We refer to our method as SoftREPA, short for Soft REPresentation Alignment. Our contributions can be summarized as follows.
@@ 2 Preliminaries
Flow Models
Suppose that we have access to samples from target distribution $X_{0}\sim q$ and source distribution $X_{1}\sim p$. The goal of flow model is to generate $X_{0}$ starting from $X_{1}$. Specifically, we define a velocity field $v_{t}({\bm{x}})$ of a flow $\psi_{t}({\bm{x}}):[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ that satisfies $\psi_{t}(X_{0})=X_{t}$ and $\psi_{1}(X_{0})=X_{1}$. Here, the $\psi_{t}$ is uniquely characterized by a flow ODE:
@@
where $v_{t}({\bm{x}}_{t}|{\bm{x}}_{0})$ defines a conditional flow $\psi_{t}({\bm{x}}_{t}|{\bm{x}}_{0})$ satisfying $\psi_{t}({\bm{x}}_{1}|{\bm{x}}_{0})={\bm{x}}_{t}$. In particular, the linear conditional flow defines the flow as ${\bm{x}}_{t}=\psi_{t}({\bm{x}}_{1}|{\bm{x}}_{0})=(1-t){\bm{x}}_{0}+t{\bm{x}}_{1}$. Then, the conditional velocity field is given by $v_{t}({\bm{x}}_{t}|{\bm{x}}_{0})=\dot{\psi}_{t}(\psi_{t}^{-1}({\bm{x}}_{t}|{\bm{x}}_{0})|{\bm{x}}_{0})={\bm{x}}_{1}-{\bm{x}}_{0}$ [23]. Thus, the conditional flow matching loss is defined as
@@
$d{\bm{x}}_{t}=\left(\mathbb{E}\left[\frac{{\bm{x}}_{t}-(1-t){\bm{x}}_{0}}{t}|{\bm{x}}_{t}\right]-\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]\right)dt=\frac{{\bm{x}}_{t}-\mathbb{E}[{\bm{x}}_{0}|{\bm{x}}_{t}]}{t}dt$ (7)
which is the probability-flow ODE (PF-ODE) of DDIM [34, 15] when ${\bm{x}}_{1}\sim p:=\mathcal{N}(0,\bm{I}_{d})$. In other words, flow models and score-based models can be used interchangeably.
Contrastive Representation Learning
Contrastive learning aims to maximize the similarity between semantically related text-image pairs while pushing apart unrelated pairs in a shared representation space [6, 31]. Formally, given a batch of $N$ image-text pairs ${(\bm{I}_{i},\bm{T}_{i})}^{N}_{i=1}$, contrastive learning optimizes a contrastive loss function, typically a variant of the InfoNCE loss:
$\mathcal{L}_{\text{CLIP}}=-\frac{1}{N}\sum^{N}_{i=1}\log\frac{\exp(\text{sim}(\bm{I}_{i},\bm{T}_{i})/\tau)}{\sum^{N}_{j=1}\exp(\text{sim}(\bm{I}_{i},\bm{T}_{j})/\tau)}$ (8)
where $\text{sim}(\bm{I}_{i},\bm{T}_{i})$ is a similarity metric such as cosine similarity or inner product between image and text embeddings, and $\tau$ denotes a temperature parameter that controls the distribution sharpness. This contrastive loss encourages image and text representations to form a joint multimodal representation space, where aligned pairs are close together, and unaligned pairs are separated. Several fundamental models have successfully applied contrastive learning for multi-modal representation learning [31, 12, 20, 21] and self-supervised learning [6].
@@
Figure 2: Network architecture and algorithmic concept of SoftREPA. (a) Learnable soft tokens of each layer are prepended to the text features across the upper layers. (b) The soft tokens are optimized to contrastively match the score with positively conditioned predicted noise while repelling the score from negatively conditioned predicted noise. This process implicitly sharpens the joint probability distribution of images and text by reducing the log probability of negatively paired conditions.
@@ 3.1 Contrastive T2I Alignment Loss
Furthermore, we define $l({\bm{x}}^{(i)},{\bm{y}}^{(j)})$ as the logit from the variation of the denoising score matching loss for the case of T2I diffusion models [40].
$l({\bm{x}},{\bm{y}})=e^{-\mathbb{E}_{t,{\bm{\epsilon}}}[\|{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t,{\bm{y}})-{\bm{\epsilon}}\|^{2}/\tau(t)]}$ (10)
@@
Here, $\tau(t)$ represents time scheduling parameter. One might consider using the denoising score matching loss, $-\mathbb{E}_{t,{\bm{\epsilon}}}[\|{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t,{\bm{y}})-{\bm{\epsilon}}\|^{2}]$, as a similarity measure instead of relying on its logit value. However, our experiments revealed that this formulation can introduce instability during training due to the unbounded nature of the similarity measure. To address this, SoftREPA introduce an exponential function to constrain the logit values, effectively stabilizing the training process.
Intuitively, as illustrated in Fig. 2(b), training a diffusion model with this loss encourage the predicted noise to align with the true noise when conditioned on a matching text description. Conversely, when conditioned on a mismatched text, the predicted noise deviates further from the true noise. This sharpens the conditional probability distribution of the trained model and increases the distinction between different text conditions, which can enhance image-text alignment in generated images.
@@ 3.2 The Soft Tokens
To distill the contrastive score matching objective above, we introduce a learnable soft token ${\bm{s}}$, which is the only trainable parameter while pretrained model remains frozen. In particular, as illustrated in fig. 2, we introduce a set of learnable soft tokens that vary between layers and are indexed by time $t$. The soft token at layer $k$ is defined using an embedding function with the number of soft tokens as $m$:
$\mathbf{s}^{(k,t)}=\text{Embedding($k,t$)}\in\mathbb{R}^{m\times d}$ (12)
At the layer $k$, the text representation is updated by concatenating the time-indexed soft token with text hidden representation from the previous layer: $\hat{\bm{H}}^{(k-1,t)}_{\text{text}}=[\mathbf{s}^{(k,t)};\bm{H}^{(k-1,t)}_{\text{text}}]\in\mathbb{R}^{(m+n)\times d}$. The soft token is prepended on text features across the layers, leading to a modified denoising function: $v_{\theta}({\bm{x}}_{t},t,{\bm{y}},{\bm{s}})$. By adopting the soft tokens, we can adjust the image and text representations to get better semantic alignment. Additionally, for computational efficiency, we approximate the expectation within $l({\bm{x}},{\bm{y}})$ using a single Monte Carlo sample, effectively replacing it as
$\tilde{l}({\bm{x}},{\bm{y}},{\bm{s}})=e^{-\|v_{\theta}({\bm{x}}_{t},t,{\bm{y}},{\bm{s}})-({\bm{\epsilon}}-{\bm{x}}_{0})\|^{2}/\tau(t)}$ (13)
@@
$\mathcal{L}_{\text{SoftREPA}}({\bm{s}})=-\mathbb{E}_{\begin{subarray}{c}({\bm{x}},{\bm{y}})\sim p_{\text{data}},t\sim U(0,1),{\bm{\epsilon}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})\end{subarray}}\log\left(\frac{\exp(\tilde{l}({\bm{x}},{\bm{y}},{\bm{s}}))}{\sum_{j}\exp(\tilde{l}({\bm{x}},{\bm{y}}^{(j)},{\bm{s}}))}\right)$ (14)
Then, the learnable token ${\bm{s}}$ is optimized by minimizing $\mathcal{L}_{\text{SoftREPA}}({\bm{s}})$. During inference, the trained soft tokens are leveraged with fixed text representation from the text encoder across layers and timesteps. The detailed procedure during the image generation process based on MM-DiT architecture is described in algorithm 1 from appendix B.
@@ 3.3 Relationship with Mutual Information
This formulation closely resembles a cross-entropy loss, where the conditional log-likelihood $\hat{l}({\bm{x}},{\bm{y}})$ can be interpreted as logits. Now, from the fact that mutual information is the expectation of the pointwise mutual information, we have
@@
where ${\bm{x}}^{(i)}$ represents the $i$-th sample in the data, $n$ denotes the number of samples in the data, and $D$ indicates the constant. Notably, Eq. (18) is very similar in form to the contrastive objective in Eq. (9), where the logit is the negative of the diffusion loss. This implies that minimizing our contrastive learning objective is closely related to maximizing the mutual information between image-text pairs under the diffusion model.
COCO val5K Efficiency
Human Preference Text Alignment Image Quality PM (GB) Latency (sec/image)
[…]
SD1.5 17.72 21.47 26.4 25.08 24.59 43.80 2.621 1.526
SD1.5 +Ours 32.89 21.50 27.33 25.18 23.43 43.38 2.621 1.547
SDXL 75.06 22.38 26.76 27.35 24.69 42.05 10.45 4.059
SDXL +Ours 85.29 22.62 26.80 28.30 26.04 42.39 10.45 4.060
SD3 94.27 22.54 26.30 28.09 31.59 42.43 18.77 4.637
SD3 +Ours 108.5 22.55 26.91 28.91 36.21 42.88 18.77 4.695
GenEval
[…]
SD3 0.68 0.99 0.86 0.56 0.85 0.27 0.55
RankDPO [16] 0.74 1.00 0.90 0.72 0.87 0.31 0.66
SD3 + Ours 0.70 1.00 0.95 0.29 0.92 0.34 0.68
Table 1: Quantitative evaluation of T2I generation with learnable soft tokens on SD1.5, SDXL, and SD3. Generation quality is evaluated on the COCO-val 5K [22] and GenEval [9] benchmark. Peak GPU memory(PM) usage and per-image latency are measured to assess computational efficiency. ImageReward, CLIP, HPS, and LPIPS are scaled by ×$\times 10^{2}$.
Human Preference Text Alignmnet Structure Background Preservation
[…]
ddim PnP 32.29 21.53 22.56 25.59 26.11 27.32 22.31 14.48 79.58
ddim PnP (Ours) 36.78 21.59 22.70 25.78 26.35 27.43 22.27 14.55 79.36
direct PnP 40.85 21.65 22.68 25.64 26.60 23.33 22.46 13.44 80.22
direct PnP (Ours) 43.32 21.70 22.80 25.83 26.76 23.41 22.40 13.53 79.93
ddim MasaCtrl -13.94 21.03 21.20 24.18 23.59 27.63 22.31 14.59 80.41
ddim MasaCtrl (Ours) -12.76 21.05 21.27 24.44 23.69 27.36 22.27 14.5 80.27
direct MasaCtrl 4.77 21.39 21.47 24.54 24.92 23.61 22.82 12.21 82.02
direct MasaCtrl (Ours) 4.48 21.40 21.49 24.69 24.93 23.32 22.77 12.19 81.85
- FlowEdit 87.70 22.16 23.19 26.72 28.04 25.48 24.37 12.55 88.58
PIEBench - FlowEdit (Ours) 102.24 22.31 23.60 27.19 28.58 24.07 24.99 12.60 88.73
ddim PnP -9.72 21.10 - 26.07 24.76 32.39 - 18.89 -
ddim PnP (Ours) -7.75 21.15 - 26.18 24.88 31.95 - 18.88 -
direct PnP -6.86 21.17 - 25.96 25.09 29.15 - 17.96 -
direct PnP (Ours) -5.56 21.22 - 26.15 25.17 28.82 - 17.99 -
ddim MasaCtrl -60.30 20.55 - 23.37 21.44 29.70 - 17.40 -
ddim MasaCtrl (Ours) -57.41 20.59 - 23.46 21.64 28.79 - 17.03 -
direct MasaCtrl -42.33 20.80 - 23.65 22.7 71.16 - 31.41 -
direct MasaCtrl (Ours) -41.20 20.81 - 23.68 22.77 71.11 - 31.26 -
- FlowEdit 38.08 21.74 - 26.05 25.53 39.68 - 15.48 -
DIV2K - FlowEdit (Ours) 46.68 21.88 - 26.39 26.03 35.66 - 14.98 -
- FlowEdit 93.70 20.72 - 22.51 26.68 33.41 - 19.92 -
C2D - FlowEdit (Ours) 114.4 21.22 - 23.43 28.35 34.05 - 20.47 -
Table 2: Quantitative evaluation of image editing performance with the use of soft tokens on PnP [37], MasaCtrl [4], Direct inversion [14], and FlowEdit [19]. FlowEdit is based on SD3, and others are based on SD1.5 architecture. The editing methods are evaluated on PIEBench [14], DIV2K [1], and Cat2Dog(C2D). ImageReward, CLIP, HPS, LPIPS, and SSIM are scaled by ×$\times 10^{2}$ and Distance is scaled by ×$\times 10^{3}$.